OpenStudy (anonymous):

express z=-4+4i in polar form please?

OpenStudy (anonymous):

\[r=\sqrt{(-4)^2+(4)^2}=4\sqrt{2}\]

OpenStudy (anonymous):

\[\theta=\tan^\left(-1\right)\left(\frac{4}{-4}\right)=-45~degrees\]

OpenStudy (anonymous):

Ty @Orian1213 , but i need the trigonometric form ..like this : z1=r(cos t + i sin t) ....

OpenStudy (anonymous):

\[z_1=4\sqrt{2}\left(\cos(-45)+i\sin(-45)\right)\]

OpenStudy (anonymous):

we can reduce this further to \[z_1=4\sqrt{2}(\cos(45)-i\sin(45))\]or simply \[z_1=4\sqrt{2}~cis~(-45)~in~short~form...\]

OpenStudy (perl):